# Real-Time Image Segmentation via Hybrid Convolutional-Transformer Architecture Search

### v2

@@ 4.2 Setup
In the search process, the initial population $P_{0}$ is set to 40 and generated by uniform random sampling. The probabilities of crossover and mutation operations are set at 0.9 and 0.02, respectively. Before the inference of a sub-network, the statistics of all Batch Normalization operations are recalculated on a random subset of training data (1000 images in Cityscapes). The number of generations is 20, and we search sub-networks with different branch numbers separately. Hence, a total of 2400 network sub-networks are searched.
@@
Table 3: Comparison of semantic segmentation on the Cityscapes. $\dagger$: speed on NVIDIA Xavier, $\sharp$: speed with TensorRT acceleration, $\star$: trained from scratch.
Methods Input Size mIoU (%) $\uparrow$ mIoU (%) $\uparrow$ FLOPs $\downarrow$ Params $\downarrow$ Speed $\uparrow$
val test (G) (M) (FPS)
ICNet [15] 1024x2048 67.7 69.5 - - 37.7
BiSeNet [16] 768x1536 69.0 68.4 - - 105.8
Fast-SCNN [17] 1024x2048 68.6 68.0 - 11.1 123.5
DF1-Seg-d8 [36] 1024x2048 72.4 71.4 - - 136.9
[…]
CAS [23] 768x1536 71.6 70.5 - - 108.0
MobilenetV3 [18] - 72.4 72.6 9.7 1.5 $10.8^{\dagger}$
SqueezeNAS [37] - 73.6 72.5 19.6 1.9 $10.2^{\dagger}$
FasterSeg [7] 1024x2048 73.1 71.5 28.2 4.4 $149^{\sharp}$
SeaFormer-S [19] 1024x2048 70.7 71.0 2.0 - 7.7
SeaFormer-B [19] 1024x2048 72.2 72.5 3.4 - 4.8
[…]
$\text{Dynamic}_{L32}$ [24] 1024x2048 79.2 80.0 242.3 - -
$\text{SETR}_{PUP}$ [12] 768x768 79.3 - 318.3 - -
[…]
$\text{OCR}_{HRNetV2-W48}$ [28] 1024x2048 80.6 - 617.2 63.8 4.6
[…]
@@ 4.4 Comparison to SOTA Models
Table 4: Comparisons of panoptic segmentation on Cityscapes validation set. $\dagger$: speed on Tesla V100-SXM2 GPU, $\ddagger$: test with multi-scale and flipping augmentation, $\star$: trained from scratch.
Methods Input Size PQ AP mIoU
$\text{TASCNet}^{\ddagger}$ [38] - 60.4 39.1 78.7
Panoptic FPN [31] - 58.1 33.0 75.7
AUNet [39] - 59.0 34.4 75.6
UPSNet [40] 1024x2048 59.3 33.3 75.2
$\text{UPSNet}^{\ddagger}$ [40] - 60.1 33.3 76.8
Seamless [41] 1025x2049 60.3 33.6 77.5
AdaptIS [42] 1024x2048 62.0 36.3 79.2
DeeperLab [43] 1025x2049 56.5 - -
$\text{SSAP}^{\ddagger}$ [44] - 61.1 37.3 -
@@
[…]
We compare our method with state-of-the-art methods on two popular benchmarks.
Cityscapes. As shown in Table 3, we can see that when the speed is more than 100 FPS, our model establishes new state-of-the-arts. HyCTAS-M consistently outperforms recent BiSeNet [16], Fast-SCNN [17], and DF1-Seg-d8 [36]. In particular, our HyCTAS-M achieves 77.1% mIoU on the Cityscapes validation set with 29.4G FLOPs and 2.4M parameters, surpassing HRNetV2-W18-Small-v1 by 6.8% with a similar model size. On the Cityscapes test set, our HyCTAS-M surpasses FasterSeg [7] by 4.2%. Moreover, the speed of our model reaches an astonishing 87.8 FPS without TensorRT acceleration. Meanwhile, our HyCTAS-S has the smallest FLOPs and parameters and achieves the fastest inference speed at 159.2 FPS. This proves that our models are very competitive when deployed on real hardware.
@@
We also evaluate HyCTAS-L on the Cityscapes validation set with the original image resolution of $1024\times 2048$. In Table  4, we see the superior PQ and speed of our HyCTAS-L. Without any inference tricks, our HyCTAS-L achieves 63.3% PQ and 81.3% mIoU, which is comparable to Panoptic-Deeplab with the HRNet-W48 [5] backbone. However, our HyCTAS-L runs twice as fast as Panoptic-Deeplab with the HRNet-W48 backbone. When we use inference data augmentation, our method achieved the best results. Specifically, our HyCTAS-L surpasses Panoptic-Deeplab [35] by 0.7% on PQ and 0.3% on mIoU. These results are achieved with only Cityscapes fine-annotated images, without using any extra data (coarse-annotated images, COCO, ImageNet, etc.).
Table 5: Comparison of semantic segmentation on the ADE20K.
Methods Encoder FLOPs Params Speed mIoU
@@
FCN [13] MobileNetV2 39.6 9.8 64.4 19.7
PSPNet [45] MobileNetV2 52.9 13.7 57.7 29.6
DeepLabV3+ [46] MobileNetV2 69.4 15.4 43.1 34.0
SegFormer [47] MiT-B0 8.4 3.8 50.5 37.4
FCN [13] ResNet-101 275.7 68.6 14.8 41.4
EncNet [48] ResNet-101 218.8 55.1 14.9 44.7
PSPNet [45] ResNet-101 256.4 68.1 15.3 44.4
DeepLabV3+ [46] ResNet-101 255.1 62.7 14.1 44.1
Auto-DeepLab [6] NAS-F48-ASPP - - - 44.0
HyCTAS(Ours) HyCTAS-L 72.7 43.2 71.8 44.9
@@
COCO. We employ the recent bottom-up panoptic segmentation framework [35] as our baseline. To ensure fast running speed, we remove the heavy context module and decoder module in Panoptic-Deeplab, which are built based on Atrous Spatial Pyramid Pooling (ASPP) [14]. We further transfer the discovered architecture, i.e., HyCTAS-L, from Cityscapes to train on COCO. To ensure the inference speed, we set the test size to $640\times 640$. Table  6 reveals that without sacrificing much accuracy, our HyCTAS-L achieves an FPS of 63.9. This high speed is over 90% faster than the closest competitor in FPS [43]. Meanwhile, we surpass Deeperlab-MNV2 [43] by a large margin, i.e., 7.5% PQ. Compared to Panoptic-Deeplab-MNV3 [35], our HyCTAS-L surpasses it on multiple performance points with 5.9% PQ. When using multi-scale and flipping, we achieve the best PQ at $640\times 640$ resolution, which is 1.6% better than Panoptic-Deeplab-ResNet-50 and 3.0% better than Deeperlab-Xception-71  [43].
Table 6: Comparison of panoptic segmentation on the COCO val2017 set. $\dagger$ denotes the speed on Tesla V100-SXM2 GPU, and $\ddagger$ denotes test with multi-scale and flipping augmentation.
Methods Input Size Speed PQ $\text{PQ}^{\text{Th}}$ $\text{PQ}^{\text{St}}$
@@
$\text{Deeperlab}_{MNV2}$ [43] 641x641 $17.2^{\dagger}$ 27.9 - -
$\text{Deeperlab}_{Xception-71}$ [43] 641x641 $10.6^{\dagger}$ 33.8 - -
[…]
HyCTAS-L 640x640 63.9 35.4 36.8 33.1
[…]
4.5 Ablation
Compared to Random Sampling. To verify the effectiveness of our searching strategy, we compared our discovered architectures with randomly sampled architectures on the Cityscapes validation set. We randomly sampled 5 architectures with 1-3 branches each from the search space. The results are presented in Table 7, in which each value is the average result of 5 random architectures. The discovered one-branch architecture, Searched-1, outperforms the random one-branch architecture by 4.3% with a smaller model size. The discovered Searched-2 and Searched-3 both surpass random architectures by more than 3%.
Table 7: Ablation study on the randomly sampled architectures on the Cityscapes validation set.
@@ 4.5 Ablation
Correlation Analysis. We sampled 10 subnetworks, training each from scratch for 40,000 iterations. The results, shown in Table 1, reveal a high correlation between the mIoU scores within the supernetwork and after retraining, with a Kendall’s $\tau$ of 0.58 and a Pearson Correlation Coefficient ($\rho$) of 0.79. This strong correlation confirms our ability to identify high-performing subnetworks within the supernetwork. Moreover, our search method not only reduces search and GPU memory costs compared to state-of-the-art methods but also efficiently yields multiple viable architectures from a single search process.
Figure 5: Visualization on COCO and Cityscapes validation set. The first row corresponds to COCO original images and the second row presents their semantic segmentation results. The third row corresponds to Cityscapes original images and the fourth row presents their semantic segmentation results. Best viewed in color.
@@ 4.6 Visualization
We visualize our predictions in Figure 5. It is obvious to see that our predictions are accurate on fine-grained details thanks to our high resolution representation learning with attention.
@@ 5 Limitations
    Sensitivity to Noise and Occlusions. Our method, like most deep architectures, can degrade in highly noisy or heavily occluded scenarios. Future work could integrate robust loss functions or data augmentation targeted at severe occlusions.
@@ 6 Conclusion
In this paper, we propose a novel HyCTAS framework to search for efficient networks with high-resolution representation and attention. We introduce a lightweight convolution module to reduce computation costs while preserving high-resolution information and a memory-efficient self-attention module to capture long-range dependencies. By combining a multi-branch search space and an efficient genetic search algorithm with multiple objectives, our searched architectures not only fuse features from selective branches for efficiency but also maintain high-resolution representations for performance. Additionally, the discovered architecture provides insights into how to use self-attention when designing new networks. Extensive experiments demonstrate that HyCTAS outperforms previous methods in both semantic segmentation and panoptic segmentation tasks while significantly improving inference speed. Future work will focus on applying this approach to more challenging tasks, such as key-point localization and object detection.
Acknowledgement
This work is jointly supported by National Key Research and Development Program of China (2016YFB1001000), National Natural Science Foundation of China (61525306, 61633021, 61721004, 61420106015, 61806194), Capital Science and Technology Leading Talent Training Project (Z181100006318030), and Beijing Science and Technology Project (Z181100008918010).
@@ Acknowledgement
H. Yu conceived the initial idea, designed the supernet, and conducted the majority of experiments. C. Wan and X. Dai contributed to the multi-objective search algorithm implementation and code maintenance. M. Liu, D. Chen, and B. Xiao provided guidance on memory-efficient self-attention design and revised the manuscript. Y. Huang, Y. Lu, and L. Wang brought expertise in high-resolution networks, coordinated the experimental setup, and offered theoretical insights on multi-branch architectures. All authors reviewed the paper, discussed experiments, and approved the final manuscript.

### v3

@@ 4.1 Benchmark and Evaluation Metrics
Figure 4: An illustration of our searched architectures with multiple branches. Green denotes lightweight convolution, red denotes memory-efficient self-attention, and gray indicates skip/fusion nodes.
@@ 4.2 Setup
In the search process, the initial population $P_{0}$ is set to 40 and generated by uniform random sampling. The crossover probability is set to 0.9, and mutation uses bit-flip over categorical genes with a per-variable rate $1/n_{\text{var}}$. Before the inference of a sub-network, the statistics of all Batch Normalization operations are recalculated on a random subset of training data (1000 images in Cityscapes). The number of generations is 20, and we search sub-networks with different branch numbers separately. Hence, a total of 2400 network sub-networks are searched.
[…]
@@
Baseline reproduction and pretraining. We re-ran from scratch (no ImageNet pretraining) a subset of baselines under our protocol and hardware, including FCN/PSPNet/DeepLabV3+ (MobileNetV2) and HRNetV2-W18S. For methods commonly reported with ImageNet pretraining (e.g., DeepLabV3+ with ResNet-101, SegFormer), we cite their original numbers and mark them as borrowed where applicable; we do not mix pretraining regimes in our claims and explicitly state the training regime for HyCTAS (from scratch) in all tables/sections.
@@ 4.3 Model Exploration
To avoid confusion, we refer to the plotted envelope as the approximate Pareto front (non-dominated set), estimated from the finite set of candidates evaluated during search.
@@ 4.4 Comparison to SOTA Models
Methods Input Size mIoU (%) $\uparrow$ FLOPs $\downarrow$ Params $\downarrow$ Speed $\uparrow$ Pretrain
val test (G) (M) (FPS)
ICNet [12] 1024x2048 67.7 69.5 - - 37.7 -
BiSeNet [13] 768x1536 69.0 68.4 - - 105.8 Y
Fast-SCNN [14] 1024x2048 68.6 68.0 - 11.1 123.5 N
DF1-Seg-d8 [41] 1024x2048 72.4 71.4 - - 136.9 Y
CAS [26] 768x1536 71.6 70.5 - - 108.0 Y
MobilenetV3 [15] - 72.4 72.6 9.7 1.5 $10.8^{\dagger}$ Y
SqueezeNAS [42] - 73.6 72.5 19.6 1.9 $10.2^{\dagger}$ Y
FasterSeg [28] 1024x2048 73.1 71.5 28.2 4.4 $149^{\sharp}$ -
SeaFormer-S [16] 1024x2048 70.7 71.0 2.0 - 7.7 Y
SeaFormer-B [16] 1024x2048 72.2 72.5 3.4 - 4.8 Y
$\text{Dynamic}_{L32}$ [29] 1024x2048 79.2 80.0 242.3 - - Y
$\text{SETR}_{PUP}$ [9] 768x768 79.3 - 318.3 - - Y
$\text{Auto-DeepLab}_{L}^{\star}$ [27] 1025x2049 80.3 80.4 695.0 44.4 0.5 N
$\text{HRNetV2}_{W18S}$ [4] 1024x2048 70.3 - 31.1 1.5 - Y
$\text{HRNetV2}_{W40}$ [4] 1024x2048 80.2 80.5 493.2 45.2 - Y
$\text{OCR}_{HRNetV2-W48}$ [33] 1024x2048 80.6 - 617.2 63.8 4.6 Y
HSNet-L [21] 512x1024 80.0 - 117.7 33.0 51.3 -
RTLinearFormer [19] 1024x2048 78.4 - 117.7 24.0 66.7 Y
$\text{HyCTAS-S}^{\star}$ 1024x2048 72.3 73.0 13.0 1.5 159.2 N
$\text{HyCTAS-M}^{\star}$ 1024x2048 77.1 75.7 29.4 2.4 87.8 N
$\text{HyCTAS-L}^{\star}$ 1024x2048 80.9 79.4 527.0 43.2 12.4 N
@@
Table 3: Comparisons of panoptic segmentation on Cityscapes validation set. $\dagger$: speed on Tesla V100-SXM2 GPU, $\ddagger$: multi-scale + flip evaluation, $\star$: trained from scratch. For entries marked with $\ddagger$, the single fixed input size is not applicable; we denote Input Size as ”MS+Flip (var)” to avoid confusion. “-” denotes values not reported by the original papers. Pretrain indicates ImageNet pretraining status (Y/N/-).
Methods Input Size PQ AP mIoU Pretrain
$\text{TASCNet}^{\ddagger}$ [43] - 60.4 39.1 78.7 Y
Panoptic FPN [36] - 58.1 33.0 75.7 Y
AUNet [44] - 59.0 34.4 75.6 Y
UPSNet [45] 1024x2048 59.3 33.3 75.2 Y
$\text{UPSNet}^{\ddagger}$ [45] - 60.1 33.3 76.8 Y
Seamless [46] 1025x2049 60.3 33.6 77.5 Y
AdaptIS [47] 1024x2048 62.0 36.3 79.2 Y
DeeperLab [48] 1025x2049 56.5 - - Y
$\text{SSAP}^{\ddagger}$ [49] - 61.1 37.3 - Y
$\text{Panoptic-DeepLab}_{HRNet-W48}$ 1025x2049 63.3 35.9 80.3 Y
$\text{Panoptic-DeepLab}_{Xception-71}$ 1025x2049 63.0 35.3 80.5 Y
$\text{HyCTAS-L}^{\star}$ 1024x2048 63.3 34.0 81.3 N
$\text{Panoptic-DeepLab}_{Xception-71}^{\ddagger}$ MS+Flip (var) 64.1 38.5 81.5 Y
$\text{HyCTAS-L}^{\ddagger\star}$ MS+Flip (var) 64.8 36.8 81.8 N
We compare our method with state-of-the-art methods on two popular benchmarks. We can see that when the speed is more than 100 FPS, our model establishes new state-of-the-arts. HyCTAS-M consistently outperforms recent BiSeNet [13], Fast-SCNN [14], and DF1-Seg-d8 [41]. In particular, our HyCTAS-M achieves 77.1% mIoU on the Cityscapes validation set with 29.4G FLOPs and 2.4M parameters, surpassing HRNetV2-W18-Small-v1 by 6.8% with a similar model size. On the Cityscapes test set, our HyCTAS-M surpasses FasterSeg [28] by 4.2%. Moreover, the speed of our model reaches an astonishing 87.8 FPS without TensorRT acceleration. Meanwhile, our HyCTAS-S has the smallest FLOPs and parameters and achieves the fastest inference speed at 159.2 FPS. This proves that our models are very competitive when deployed on real hardware.
@@
Our HyCTAS-L achieves superior performance compared to recent transformer-based methods such as HSNet-L [21], which uses hierarchical similarity networks for few-shot segmentation, and RTLinearFormer [19], which employs linear attention mechanisms for real-time performance. These comparisons demonstrate that our search-based approach can discover more effective hybrid architectures than handcrafted designs.
@@
We also evaluate HyCTAS-L on the Cityscapes validation set with the original image resolution of $1024\times 2048$. In Table  3, we see the superior PQ and speed of our HyCTAS-L. Without any inference tricks, our HyCTAS-L achieves 63.3% PQ and 81.3% mIoU, which is comparable to Panoptic-Deeplab with the HRNet-W48 [4] backbone. However, our HyCTAS-L runs twice as fast as Panoptic-Deeplab with the HRNet-W48 backbone. When we use inference data augmentation, our method achieved the best results. Specifically, our HyCTAS-L surpasses Panoptic-Deeplab [40] by 0.7% on PQ and 0.3% on mIoU. These results are achieved with only Cityscapes fine-annotated images, without using any extra data (coarse-annotated images, COCO, ImageNet, etc.).
Table 4: Comparison of semantic segmentation on ADE20K. “-” denotes values not reported by the original papers. Pretrain indicates ImageNet pretraining status (Y/N/-).
Methods Encoder FLOPs Params Speed mIoU Pretrain
@@
FCN [10] MobileNetV2 39.6 9.8 64.4 19.7 Y
PSPNet [50] MobileNetV2 52.9 13.7 57.7 29.6 Y
DeepLabV3+ [51] MobileNetV2 69.4 15.4 43.1 34.0 Y
SegFormer [18] MiT-B0 8.4 3.8 50.5 37.4 Y
FCN [10] ResNet-101 275.7 68.6 14.8 41.4 Y
EncNet [52] ResNet-101 218.8 55.1 14.9 44.7 Y
PSPNet [50] ResNet-101 256.4 68.1 15.3 44.4 Y
DeepLabV3+ [51] ResNet-101 255.1 62.7 14.1 44.1 Y
Auto-DeepLab [27] NAS-F48-ASPP - - - 44.0 N
UperNet [53] DeiT-S [54] - 43.0 - 41.0 Y
UperNet [53] Vim-S [23] - 46.0 - 44.9 Y
AFFormer-B [22] - - 3.0 49.6 41.8 Y
RTLinearFormer [19] - 14.7 24.0 107.3 35.7 Y
HyCTAS-L HyCTAS-L 72.7 43.2 71.8 44.9 N
@@
Our approach shows improved results over recent efficient transformer models including UperNet with DeiT-S [54] and Vim-S [23] backbones, as well as AFFormer-B [22]. These comparisons highlight the effectiveness of our search strategy in discovering optimal hybrid architectures that balance accuracy and efficiency better than handcrafted transformer-based designs.
COCO. We employ the recent bottom-up panoptic segmentation framework [40] as our baseline. To ensure fast running speed, we remove the heavy context module and decoder module in Panoptic-Deeplab, which are built based on Atrous Spatial Pyramid Pooling (ASPP) [11]. We further transfer the discovered architecture, i.e., HyCTAS-L, from Cityscapes to train on COCO. To ensure the inference speed, we set the test size to $640\times 640$. Table  5 reveals that without sacrificing much accuracy, our HyCTAS-L achieves an FPS of 63.9. This high speed is over 90% faster than the closest competitor in FPS [48]. Meanwhile, we surpass Deeperlab-MNV2 [48] by a large margin, i.e., 7.5% PQ. Compared to Panoptic-Deeplab-MNV3 [40], our HyCTAS-L surpasses it on multiple performance points with 5.9% PQ. When using multi-scale and flipping, we achieve the best PQ at $640\times 640$ resolution, which is 1.6% better than Panoptic-Deeplab-ResNet-50 and 3.0% better than Deeperlab-Xception-71  [48].
Our HyCTAS-L demonstrates competitive advantages over recent methods such as AFFormer-B [22] and RTLinearFormer [19] on the COCO dataset, demonstrating the generalization capability of our searched architectures across different benchmarks. These results further validate that our search-based approach can discover more effective hybrid designs than handcrafted transformer-based models.
Table 5: Comparison of panoptic segmentation on the COCO val2017 set. $\dagger$ denotes speed on Tesla V100-SXM2 GPU; $\ddagger$ denotes multi-scale + flip evaluation. Entries without $\dagger$ are measured by us on RTX 3090 at $\times640$, batch size 1. For $\ddagger$ entries, a single fixed input size and FPS are not applicable; we denote Input Size as ”MS+Flip (var)” and FPS as ”n/a” for clarity. “-” denotes values not reported by the original papers. Pretrain indicates ImageNet pretraining status (Y/N/-).
Methods Input Size Speed PQ $\text{PQ}^{\text{Th}}$ $\text{PQ}^{\text{St}}$ Pretrain
@@
$\text{Deeperlab}_{\text{MNV2}}$ [48] 641x641 $17.2^{\dagger}$ 27.9 - - Y
$\text{Deeperlab}_{\text{Xception-71}}$ [48] 641x641 $10.6^{\dagger}$ 33.8 - - Y
$\text{Panoptic-DeepLab}_{\text{MNV3}}$ [40] 641x641 $13.5^{\dagger}$ 29.5 - - Y
$\text{Panoptic-DeepLab}_{\text{ResNet-50}}$ [40] 641x641 $6.0^{\dagger}$ 35.2 - - Y
AFFormer-B [22] 640x640 46.5 35.1 - - Y
RTLinearFormer [19] 640x640 150.2 31.1 - - Y
HyCTAS-L 640x640 63.9 35.4 36.8 33.1 N
$\text{HyCTAS-L}^{\ddagger}$ MS+Flip (var) n/a 36.8 38.5 34.5 N
Table 6: Ablation study on the randomly sampled, locally searched, and NSGA-II searched architectures on the Cityscapes validation set.
@@
4.5 Ablation
[…]
Selection of Search-1/2/3. From the approximate Pareto fronts for each branch configuration (1-, 2-, and 3-path), we manually select Search-1/2/3 by jointly considering FLOPs, parameter count, and latency; these selections correspond to HyCTAS-S/M/L. More broadly, our method yields a family of models along these fronts, allowing practitioners to choose configurations that match their deployment priorities (e.g., maximizing mIoU for accuracy or minimizing latency for speed).
Selection of Local-1/2/3. We pool all evaluated Local Search candidates across seeds and iterations and compute the approximate Pareto front w.r.t. (mIoU, FLOPs). Local-1/2/3 are then chosen from this pooled front as the models whose FLOPs are closest to the already selected Searched-1/2/3.
@@ 4.5 Ablation
Correlation Analysis. We sampled 30 subnetworks, training each from scratch for 40,000 iterations. We observe a positive rank and linear correlation between mIoU measured within the supernetwork and after retraining (Kendall’s $\tau{=}0.51$, Pearson’s $\rho{=}0.70$), suggesting that supernet evaluation is a useful proxy for ranking.
Figure 5: Visualization on COCO and Cityscapes validation set. The first row corresponds to COCO original images and the second row presents their semantic segmentation results. The third row corresponds to Cityscapes original images and the fourth row presents their semantic segmentation results. Best viewed in color. Zoom in for details; predictions retain fine boundaries due to high-resolution branches.
@@ 4.6 Visualization
[…]
@@ 5 Limitations
    Sensitivity to Noise and Occlusions. Our method, like most deep architectures, may degrade under severe noise or occlusion. Future work could integrate robust loss functions or data augmentation targeted at severe occlusions.
@@
  • 4.
    Search cost and hardware dependency. Although the search is relatively efficient, the supernet pretraining plus NSGA-II still requires notable GPU hours, and the latency objective is calibrated to a specific GPU/driver stack. Architectures on the approximate Pareto front may shift if the deployment hardware changes (e.g., mobile DSP vs. server GPU).
  • 5.
    Calibration and small-batch effects. Latency and accuracy during search are measured with batch size 1 and BatchNorm re-estimation on a subset; distribution shifts at deployment (e.g., different preprocessing or batch size) can slightly change the accuracy–latency trade-off.
  • 6.
    Transferability across tasks. While we show promising panoptic results, transfer to tasks with different optimization targets (e.g., detection or pose) is not guaranteed; the search objectives would likely need retuning.
To mitigate these issues, future work will (i) incorporate portable latency proxies (e.g., per-operator cost models) for cross-hardware robustness; (ii) explore differentiable or predictor-guided multi-objective search to further reduce GPU hours; and (iii) extend the search space with domain-adaptive modules and normalization layers that are less sensitive to batch effects.
@@ 6 Conclusion
In this paper, we propose a novel HyCTAS framework to search for efficient networks with high-resolution representation and attention. We introduce a lightweight convolution module to reduce computation costs while preserving high-resolution information and a memory-efficient self-attention module to capture long-range dependencies. By combining a multi-branch search space and an evolutionary computation-based approach (NSGA-II) with multiple objectives, our searched architectures not only fuse features from selective branches for efficiency but also maintain high-resolution representations for performance. Additionally, the discovered architecture provides insights into how to use self-attention when designing new networks. Extensive experiments demonstrate that HyCTAS outperforms previous methods in both semantic segmentation and panoptic segmentation tasks while significantly improving inference speed. Future work will focus on applying this approach to more challenging tasks, such as key-point localization and object detection.
In summary, HyCTAS conducts a single multi-objective search to produce an approximate Pareto set of hybrid convolution–attention architectures that operate at high resolution and deliver strong accuracy–latency trade-offs on Cityscapes, ADE20K, and COCO, all trained from scratch. The search further reveals systematic placement patterns of attention across branches (shallower for single-branch, deeper for multi-branch) that can inform future network design.
The primary contribution is the NAS strategy and search space for multi-branch high-resolution segmentation with explicit latency as an objective; the lightweight convolution and memory-efficient attention are pragmatic operator choices that enable feasible search rather than architectural novelties.
Regarding limitations and outlook, the latency estimates are hardware-specific and the search still requires GPU hours. Future directions include portable latency models and predictor-guided search to reduce cost, analyzing attention placement under different objective weights, and extending HyCTAS to additional tasks.
@@
H. Yu conceived the initial idea, designed the supernet, and conducted the majority of experiments. C. Wan and X. Dai contributed to the multi-objective search algorithm implementation and code maintenance. M. Liu, D. Chen, and B. Xiao provided guidance on memory-efficient self-attention design and revised the manuscript. Y. Huang, Y. Lu, and L. Wang brought expertise in high-resolution networks, coordinated the experimental setup, and offered theoretical insights on multi-branch architectures. All authors reviewed the paper, discussed experiments, and approved the final manuscript. All authors agree to the author list and order and are accountable for all aspects of the work.